A fence 10 ft high runs parallel to the wall of a house at a distance of 6 ft. Find the length of the shortest ladder that exPplease show you work! I cannot figure out the steps
Answer #1

A ㅌ BE CLet DC=10ft

BC=6ft

Let CE=b

AB=a

In triangle ABE and DCE

a/6+b=10/b

a=(60+10b)/b

Let AE=L

L^2= a^2 + (b+6)^2

=((60+10b)/b)^2 +(b+6)^2

Differentiate

2L dL/db= 2(60+10b)/b(-60/b^2)+2(b+6)

=-120(60+10b)/b^3+2(b+6)

dL/dB=0

0=2(b+6)(1-600/b^3)

b=-6

b=600^(1/3)=8.434

L^2=((60+10*8.434)/8.434)^2+(8.434+6)^2

=292.897+208.34035=501.23735

L=22.3883